OpenStudy (anonymous):

Find g of f if f(x) = x + 7 and g(x) = 3x – 2. A. 3x + 5 B. 3x - 5 C. 3x + 19 D. -3x - 19

OpenStudy (anonymous):

Your answer is A. replace the x in f(x) =x +7 with (3x-2) g(f(x)=(3x-2)+7 g(f(x)=3x+5